A storage tank contains 2 moles of Ar, 3 moles of O2, and 5 moles




of N2 at a total pressure of 1000 torr. Calculate partial pressure of




each gas.





1
Expert's answer
2022-05-31T09:04:03-0400

partial pressure can be calculated as follows:

P(gas 1) = x(gas 1) * P(Total);

where x(gas 1) = no of moles(gas 1)/ no of moles(total). As you can see the above formulae does not require the individual volumes of the gases or the total volume.

1.

x(Ar) = 2mol / 10mol = 0,2

P(Ar) = x(Ar) * P(total) =0,2 * 1000 =200 torr

2.

x(O2) = 3mol / 10mol = 0,3

P(O2) = 0,3 * 1000 = 300 torr

3.

x(N2) = 5mol / 10mol = 0,5

P(N2) = 0,5 * 1000 = 500 torr

ANSWER:

Ar : 200 torr

O2 : 300 torr

N2 : 500 torr